# Newton numbers, vanishing polytopes and algebraic degrees

### v1

@@ 5.4 Algebraic degrees
We show that the formulas for the maximum likelihood degree from [CHKS06, Theorem 15] and [LNRW23], the Euclidean distance degree (introduced in [DHOST16]) from [BSW22] and [TT24], and the polar degree from [Huh13, Corollary 4.6.] are all implied by properties of the $e$-Newton number given in §5.3. Our formulas generalize those for maximum likelihood and Euclidean distance degrees, while for polar degree we provide new formulas (including expressions via the mixed volume).
We also discuss geometrically interesting cases, where the maximum likelihood degree is equal to 00 or $1$ (see [Huh14] and [Es24c, Theorem 1.11]), and the polar degree is equal to 00 or $1$ (see [Huh13, Example 4.8. and below] and [CRS08]).
The following theorem is [Es24b, Example 1.7(6)] with several minor additions and corrections. We consider all the algebraic degrees in separate subsubsections below. In the following theorem we use the usual Lagrangian multiples, and the ones that we call ‘‘outer’’ Lagrangian multiples and ‘‘dehomogenizing’’ Lagrangian multiples. We call them this way since ‘‘outer’’ Lagrangian multiples let us count the critical points outside the corresponding variety and ‘‘dehomogenizing’’ Lagrangian multiples let us count the degree of the gradient map by extracting a single point out of a projective line (for more detailed see the corresponding subsubsections below).
@@
    For subvarieties (see [Huh13] and [HS]), we calle it the ML-degree.
@@
    $\partial_{Q_{u}^{\star}}(\mathbf{\Theta}_{\text{MML}}(z,\lambda))=0\text{\quad, where}$
@@
    Eucledean distance degree (ED-degree) from [DHOST16]
@@
    Denote by $\mathbf{\Delta}_{n+1}$ the set of vertices of the unit simplex in $\mathbb{Z}^{n+1}$. In the Newton non-degenerate case with support $\mathbf{P}$ the P-degree (i.e. the number of the preimages of a generic point $\mathbf{c}\in\mathbb{C}\mathbb{P}^{n}$) is equal to:
@@
    Note that the dimension can be reduced by $1$ (see Remark 5.42).
@@ 5.4.1 Mapping maximum likelihood degree
Let us show that ‘‘outer’’ Lagrangian from Theorem 5.34(1) indeed compute the MML-degree.
@@
Assume all $u_{i}$ are non-zero (otherwise we may eliminate the corresponding $\Theta_{i}$; also note that the formulas below can be extended to cases where some $u_{i}=0$, but not all simultaneously)
The equation $\partial_{Q_{u}^{\star}}(\mathbf{\Theta}_{\text{MML}}(z,\lambda))=0$ in $(\mathbb{C}\setminus 0)^{n}$ in variables $\lambda_{1},\dots,\lambda_{m},z_{1},\dots,z_{n-m}$ is equivalent to:
$\begin{cases}\Theta_{1}\lambda_{1}+\dots+\Theta_{m}\lambda_{m}+c=0\\ \frac{\lambda_{1}\Theta_{1}}{u_{1}}=\dots=\frac{\lambda_{m}\Theta_{m}}{u_{m}}\\ \sum\limits_{i=1}^{m}\lambda_{i}z_{1}\frac{\partial\Theta_{i}}{\partial z_{1}}=0\\ \quad\dots\\ \sum\limits_{i=1}^{m}\lambda_{i}z_{n-m}\frac{\partial\Theta_{i}}{\partial z_{n-m}}=0\end{cases}$
Note that the first two equations have at most one solution in $\lambda$’s, existing if and only if all $\Theta_{i}\neq 0$. When a solutions exist, all $\lambda_{i}$ are non-zero. The remaining equations give exactly the conditions for critical point in the torus.
@@
[CHKS06, Theorem 19.] Under certain non-degeneracy conditions, the MML-degree of is equal to the signed Euler characteristic for almost all $u\subset\mathbb{Z}^{n}$:
@@
The formula for the MML-degree for Newton non-degenerate polynomials and with certain smoothness assumption from [CHKS06, Theorem 15.] is of the following form:
$\sum_{1\leq i_{1}\leq\cdots\leq i_{n-m}\leq m}MV(P_{i_{1}},\dots,P_{i_{n-m}})\,\,\,-\!\mathop{\sum_{j\in\{1,\dots,r\}}}_{1\leq i_{1}\leq\cdots\leq i_{{n-m}-1}\leq m}MV(P_{i_{1}},\dots,P_{i_{{n-m}-1}};\tau_{\{j\}})\\ +\,\,\mathop{\sum_{\{j_{1},j_{2}\}\subset\{1,\dots,r\}}}_{1\leq i_{1}\leq\cdots\leq i_{{n-m}-2}\leq m}MV(P_{i_{1}},\dots,P_{i_{{n-m}-2}};\tau_{\{j_{1},j_{2}\}})\quad+\quad\cdots\quad+\\ (-1)^{n-m}\sum_{\{j_{1},\dots,j_{n-m}\}\subset\{1,\dots,r\}}MV(\emptyset;\tau_{\{j_{1},\ldots,j_{n-m}\}}).$
@@
For generic $u$ (and after a generic shift by $\alpha\in\mathbb{Z}^{n-m}$ corresponding to the multiplication by a generic monomial $z^{\alpha}$), only the first summand left. It agrees with the computation of the corresponding Euler characteristic (Theorem 5.35) via the inclusion-exclusion principle applied to the BKK-formula Theorem 5.7. The same Euler characteristic also computes the degree of so called ‘‘toric polar map’’ (see [FMS24, Theorem 1.1]).
@@ 5.4.2 Maximum likelihood degree of very affine varieties
Theorem 5.34 (2) follows directly from standard Lagrange multipliers techniques. Recall that the ML-degree is equal to the signed Euler characteristic.
@@
$\text{MLdeg}(V_{\mathbf{f}},u)=MV(P_{1},\dots,P_{m},\underbrace{P,\dots,P}_{n-m}),$
where $P=u\ast P_{1}\ast\dots\ast P_{m}\subset\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$.
Note that Lemma 5.30 implies this example. For generic $u$, Lemma 5.32 provides a formula for the ML-degree as the (usual) Newton number. The result is consistent with the BKK-theorem (see Lemma 5.10). Note that $P$ is not convenient in the cone $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ and we do not know how to obtain a formula for the ML-degree via the $\ell$-Newton number.
@@
  • ML-degree 1 (via rational map called ‘‘Horn uniformization’’ $\mathbb{C}\mathbb{P}^{m-1}\dashrightarrow V$, see [Huh14])
@@
Question 5.40.
Can we relate:
@@
    The classification of very affine varieties with ML-degree 00 or $1$
@@
    Criteria for the mixed volume to be equal to 00 or $1$ (Propositions 5.5, 5.6)
@@ 5.4.3 Euclidean distance degree
$\text{EDdeg}(V_{\mathbf{f}})=MV(\operatorname{Conv}(\mathbf{P}^{\rho}\setminus E_{1}),\dots\operatorname{Conv}(\mathbf{P}^{\rho}\setminus E_{n})),$
@@
The $e$-Newton number formula (Theorem 5.34(3)) is more general since it does not require the supports to contain the origin. In this more general case the mixed volume and the (usual or $\ell$-) Newton number formulas are no longer correct.
@@ 5.4.4 Polar degree
Proof that the latter remark computes the preimages under the gradient map is the same as above. The equivalence to the (usual) Newton number is implied by Lemma 5.32.
Recall the Huh’s formula for the degree of the gradient map [Huh13, Corollary 4.6]. Consider a homogeneous degree $d$ polynomial $f\in\mathbb{C}[z_{0},\dots,z_{n}]$. Denote by $P\subset d\Delta^{\circ}$ the Newton polytope of $f$. Define
$m_{i}(P):=\text{MV}_{n}(\underbrace{\Delta^{\circ},\ldots,\Delta^{\circ}}_{n-i},\underbrace{P,\ldots,P}_{i})\quad\text{for}\ i=0,\ldots,n.$
@@
([Huh13, Corollary 4.6.]) Suppose that $f$ is nondegenerate with respect to its Newton polytope. Then the degree of the gradient map $\text{grad}(f):\mathbb{C}\mathbb{P}^{n}\to\mathbb{C}\mathbb{P}^{n}$ is the sum
@@
Denote by $E_{0}^{d},\dots,E_{n}^{d}$ the facets of $C_{d}$. Denote by $D_{i}$ the polytope $\operatorname{Conv}((\mathbf{P}\ast\mathbf{\Delta}^{\circ})\setminus E^{d}_{i})\subset C_{d}$. Then for generic homogeneous degree $d$ polynomials with support $\mathbf{P}\subset\mathbb{Z}^{n+1}$ we have the following formula for the polar degree:
$\text{Pdeg}(f)=MV(D_{0},\dots,D_{n})$
@@
The most interesting cases are the cases of the polar degree equal to zero (vanishing hessian hypersurfaces) and equal to one (homaloidal hypersurfaces corresponding to Cremona transformations), see e.g. [CRS08]. Note that due to the Proposition 5.44, the Propositions 5.5 and 5.6 provide criteria for these cases (in Newton non-degenerate case).
Question 5.46.
Can we classify polytopes $P\subset d\Delta^{\circ}$, so that $\nu_{C_{d}}(P\ast\Delta^{\circ})$ is equal to 00 or $1$?
For instance, we have a conjecture concerning Newton non-degenerate homaloidal hypersurfaces. It is easy to verify it for the examples [Huh13, below Example 4.9] and we did not find any counterexamples.
@@
Consider the abstract lattice polyhedral poset $\mathcal{P}^{\bullet}$ defined as the agglutination of $P\ast\Delta^{\circ}$ and the face poset of the unit $(n+1)$-simplex $\Delta^{\bullet}$ along the face poset of the $n$-simplex $\Delta^{\circ}$. If we forget about the fact that the poset $\mathcal{P}^{\bullet}$ consists of lattice polytopes, we see that it is isomorphic to the (non-lattice) subdivision of $\operatorname{Conv}(O_{d},\Delta^{\circ},P)$ into $P\ast\Delta^{\circ}$ and $\operatorname{Conv}(O_{d},\Delta^{\circ})$, where $O_{d}$ is the apex of the cone $C_{d}$. Note that $\operatorname{Conv}(O_{d},\Delta^{\circ},P)$ is a (non-lattice) convenient polytope in $C_{d}$. This induces strong formal subdivision with boundary $\sigma:(\mathcal{P}^{\bullet},\mathcal{P}^{\bullet}_{B})\to\mathcal{C}^{\pm}$, where $\mathcal{C}$ is the face poset of the cone $C_{d}$ (or, equivalently, the face poset of $\Delta^{\circ}$).
@@
$\text{Pdeg}(f)=\nu_{{C_{d}}}(P\ast\Delta^{\circ})=\frac{1}{2}\ell^{*}_{\mathcal{C}}(\mathcal{P}^{\bullet}\sharp_{\mathcal{P}^{\bullet}_{B}}\mathcal{P}^{\bullet};1)$
@@
The poset $\mathcal{C}$ is boolean, so the local $h_{\mathcal{C}}^{*}$-polynomial is the alternating sum of the corresponding $h_{\mathcal{C}}^{*}$-polynomials. The evaluation of the $h_{\mathcal{C}}^{*}$-polynomial of polytopes at $t=1$ is the lattice volume. The corresponding alternating sum of volumes for $\Delta^{\bullet}$ equals 00 and for $\Delta^{\circ}\ast P$ equals $2\nu_{{C_{d}}}(P\ast\Delta^{\circ})$. ∎
@@
$Pdeg(f)\leq\ell^{*}(P\ast\Delta^{\circ};1)$
6 $B_{k}$-polytopes

### v2

@@ 5.5 Algebraic degrees
We show that the formulas for the maximum likelihood degree from [CHKS06, Theorem 15] and [LNRW23], the Euclidean distance degree (introduced in [DHOST16]) from [BSW22] and [TT24], and the polar degree from [Huh13, Corollary 4.6] are all implied by properties of the $e$-Newton number given in §5.3. Our formulas generalize those for maximum likelihood and Euclidean distance degrees, while for polar degree we provide new formulas (including expressions via the mixed volume).
We also discuss geometrically interesting cases, where the maximum likelihood degree is equal to 0 or $1$ (see [Huh14] and [Es24c, Theorem 1.11]), and the polar degree is equal to 0 or $1$ (see [Huh13, Example 4.8. and below] and [CRS08]).
The following theorem is [Es24b, Example 1.7(6)] with several minor additions and corrections. We consider all the algebraic degrees in separate subsubsections below. In the following theorem we use the usual Lagrangian multiples, and the ones that we call ‘‘outer’’ Lagrangian multiples and ‘‘dehomogenizing’’ Lagrangian multiples. We call them this way since ‘‘outer’’ Lagrangian multipliers let us count the critical points outside the corresponding variety, and ‘‘dehomogenizing’’ Lagrangian multiples let us count the degree of the gradient map by extracting a single point from a projective line (for more details, see the corresponding subsubsections below).
@@
    For subvarieties (see [Huh13] and [HS]), we call it the ML-degree.
@@
    $\partial_{Q_{u}^{\star}}(\mathbf{\Theta}_{\text{MML}}(z,\lambda))=0\text{,\quad where}$
@@
    Euclidean distance degree (ED-degree) from [DHOST16]
@@
    Denote by $\mathbf{\Delta}_{n+1}$ the set of vertices of the unit $(n+1)$-simplex in $\mathbb{Z}^{n+1}$ (containing the origin). In the Newton non-degenerate case of homogeneous polynomial with support $\mathbf{P}\subset\mathbb{Z}^{n+1}$ the P-degree (i.e. the number of the preimages of a generic point $\mathbf{c}\in\mathbb{C}\mathbb{P}^{n}$) is equal to:
@@
    Note that the dimension $n+1$ of the simplex $\mathbf{\Delta}_{n+1}$ can be reduced to $n$ (see Remark 5.49).
@@ 5.5.1 Mapping maximum likelihood degree
Let us show that ‘‘outer’’ Lagrangian multipliers from Theorem 5.40(1) indeed compute the MML-degree.
@@
Assume all $u_{i}$ are nonzero (otherwise we may eliminate the corresponding $\Theta_{i}$; also note that the formulas below can be extended to cases where some $u_{i}=0$, but not all simultaneously). The equation $\partial_{Q_{u}^{\star}}(\mathbf{\Theta}_{\text{MML}}(z,\lambda))=0$ in $(\mathbb{C}\setminus 0)^{n}$ in variables $\lambda_{1},\dots,\lambda_{m},z_{1},\dots,z_{n-m}$ is equivalent to:
$\begin{cases}\Theta_{1}\lambda_{1}+\ldots+\Theta_{m}\lambda_{m}+c=0\\ \frac{\lambda_{1}\Theta_{1}}{u_{1}}=\ldots=\frac{\lambda_{m}\Theta_{m}}{u_{m}}\\ \sum\limits_{i=1}^{m}\lambda_{i}z_{1}\frac{\partial\Theta_{i}}{\partial z_{1}}=0\\ \quad\vdots\quad\vdots\quad\vdots\\ \sum\limits_{i=1}^{m}\lambda_{i}z_{n-m}\frac{\partial\Theta_{i}}{\partial z_{n-m}}=0\end{cases}$
Note that the first two equations have at most one solution in $\lambda$’s, existing if and only if all $\Theta_{i}\neq 0$. When a solutions exist, all $\lambda_{i}$ are nonzero. The remaining equations give exactly the conditions for critical point in the torus.
@@
[CHKS06, Theorem 19] Under certain non-degeneracy conditions, the MML-degree of is equal to the signed Euler characteristic for almost all $u\subset\mathbb{Z}^{n}$:
@@
The formula for the MML-degree for Newton non-degenerate polynomials and with certain smoothness assumptions from [CHKS06, Theorem 15] is of the following form:
$\sum_{1\leq i_{1}\leq\cdots\leq i_{n-m}\leq m}\operatorname{MV}(P_{i_{1}},\dots,P_{i_{n-m}})\,\,\,-\!\mathop{\sum_{j\in\{1,\dots,r\}}}_{1\leq i_{1}\leq\cdots\leq i_{{n-m}-1}\leq m}\operatorname{MV}(P_{i_{1}},\dots,P_{i_{{n-m}-1}};\tau_{\{j\}})\\ +\,\,\mathop{\sum_{\{j_{1},j_{2}\}\subset\{1,\dots,r\}}}_{1\leq i_{1}\leq\cdots\leq i_{{n-m}-2}\leq m}\operatorname{MV}(P_{i_{1}},\dots,P_{i_{{n-m}-2}};\tau_{\{j_{1},j_{2}\}})\quad+\quad\cdots\quad+\\ (-1)^{n-m}\sum_{\{j_{1},\dots,j_{n-m}\}\subset\{1,\dots,r\}}\operatorname{MV}(\emptyset;\tau_{\{j_{1},\ldots,j_{n-m}\}}).$
@@
For generic $u$ (and after a generic shift by $\alpha\in\mathbb{Z}^{n-m}$ corresponding to the multiplication by a generic monomial $z^{\alpha}$), only the first summand remains. It agrees with the computation of the corresponding Euler characteristic (Theorem 5.41) via the inclusion-exclusion principle applied to the BKK-formula Theorem 5.8. The same Euler characteristic also computes the degree of the so-called ‘‘toric polar map’’ (see [FMS24, Theorem 1.1]).
@@ 5.5.2 Maximum likelihood degree of very affine varieties
Theorem 5.40 (2) follows directly from standard Lagrange multipliers techniques. Recall that the ML-degree is equal to the signed Euler characteristic in the smooth case.
@@
$\text{MLdeg}(V_{\mathbf{f}},u)=\operatorname{MV}(P_{1},\dots,P_{m},\underbrace{P,\dots,P}_{n-m}),$
where $P=\{u\}\ast P_{1}\ast\dots\ast P_{m}\subset\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$.
Note that Example 5.44 follows from Lemma 5.31. For generic $u$, Lemma 5.33 provides a formula for the ML-degree as the (usual) Newton number in $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$. The result is consistent with the BKK-theorem (see Lemma 5.11). Note that $P$ is not convenient in the cone $\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ and we do not know how to obtain a formula for the ML-degree via the $\ell$-Newton number.
@@
  • ML-degree 1 (via a rational map called ‘‘Horn uniformization’’ $\mathbb{C}\mathbb{P}^{m-1}\dashrightarrow V$, see [Huh14])
@@
Question 5.46.
Can we relate the following for Newton non-degenerate complete intersections?
@@
    The classification of very affine varieties with ML-degree 0 or 11
@@
    Criteria for the mixed volume to be equal to 0 or 11 (Propositions 5.6, 5.7)
@@ 5.5.3 Euclidean distance degree
$\text{EDdeg}(V_{\mathbf{f}})=\operatorname{MV}(\operatorname{Conv}(\mathbf{P}^{\rho}\setminus E_{1}),\dots,\operatorname{Conv}(\mathbf{P}^{\rho}\setminus E_{n})),$
@@
The $e$-Newton number formula (Theorem 5.40(3)) is more general since it does not require the supports to contain the origin. In this more general case the formulas with mixed volume and the (usual or $\ell$-) Newton number are no longer correct. Here is an example.
Example 5.48.
Consider one polynomial $f\in\mathbb{C}[z_{1},z_{2},z_{3},z_{4}]$ in $4$ variables with support $\mathbf{P}$ consisting of the following $6$ points and generic coefficients:
AA $B$ $X$ $Y$ $V$ $W$
$e_{1}$ $e_{2}$ $e_{1}+e_{3}$ $e_{1}+e_{4}$ $e_{2}+e_{3}$ $e_{2}+e_{4}$
Figure 4: The support $\mathbf{P}$ in dimension 44
Denote by $\rho$ a generic quadratic distance function in $4$ variables. Let us verify that the system of derivatives
$\frac{\partial(\lambda f+\rho)}{\partial z_{1}}=\frac{\partial(\lambda f+\rho)}{\partial z_{2}}=\frac{\partial(\lambda f+\rho)}{\partial z_{3}}=\frac{\partial(\lambda f+\rho)}{\partial z_{4}}=\frac{\partial(\lambda f+\rho)}{\partial z_{\lambda}}=0$ (6)
with the supports
$(\mathbf{\Delta}_{\rho}\ast\mathbf{P})\setminus E_{1},\quad(\mathbf{\Delta}_{\rho}\ast\mathbf{P})\setminus E_{2},\quad(\mathbf{\Delta}_{\rho}\ast\mathbf{P})\setminus E_{3},\quad(\mathbf{\Delta}_{\rho}\ast\mathbf{P})\setminus E_{4},\quad\mathbf{P}$
is degenerate for the covector $\xi=(-1,-1,0,0,1)$ in the Bernstein-Kouchnirenko sense Theorem 5.4. The affine span of the support face $(\mathbf{\Delta}_{\rho}\ast\mathbf{P})^{\xi}$ has dimension $4$ and contains the origin. The lattice $\text{aff}(\mathbf{\Delta}_{\rho}\ast\mathbf{P})^{\xi}\cap\mathbb{Z}^{5}$ has the lattice basis
$v_{1}=(1,0,0,0,1)\quad v_{2}=(0,1,0,0,1)\quad v_{3}=(0,0,1,0,0)\quad v_{4}=(0,0,0,1,0).$
In this basis the face $(\mathbf{\Delta}_{\rho}\ast\mathbf{P})^{\xi}$ consists of the following points:
$O$ $A^{\prime}$ $B^{\prime}$ $X^{\prime}$ $Y^{\prime}$ $V^{\prime}$ $W^{\prime}$ $R_{1}$ $S_{1}$ $R_{2}$ $S_{2}$
0 $v_{1}$ $v_{2}$ $v_{1}+v_{3}$ $v_{1}+v_{4}$ $v_{2}+v_{3}$ $v_{2}+v_{4}$ $v_{3}$ $v_{4}$ $2v_{3}$ $2v_{4}$
Degeneracy of the system (6) for $\xi$ is equivalent to the inequality $\nu^{e}_{O}((\mathbf{\Delta}_{\rho}\ast\mathbf{P})^{\xi})>0$ in the lattice span by $v_{1},v_{2},v_{3},v_{4}$. By Lemma 5.31 and Proposition 5.7 we have:
$\nu^{e}_{O}((\mathbf{\Delta}_{\rho}\ast\mathbf{P})^{\xi})=\operatorname{MV}(A^{\prime}X^{\prime}Y^{\prime},B^{\prime}V^{\prime}W^{\prime},X^{\prime}V^{\prime}R_{1}R_{2},Y^{\prime}W^{\prime}R_{2}S_{2})=1,$
thus the system (6) is degenerate for $\xi$.
Using SageMath we computed that:
$\operatorname{MV}(\operatorname{Conv}((\mathbf{\Delta}_{\rho}\ast\mathbf{P})\setminus E_{1}),\dots,\operatorname{Conv}((\mathbf{\Delta}_{\rho}\ast\mathbf{P})\setminus E_{4}),\operatorname{Conv}(\mathbf{P}))=7,$
and that the number of solutions of the system (6) in $\mathbb{C}^{5}$ is equal to $6$. So our SageMath computations also confirm that the system (6) is degenerate.
@@ 5.5.4 Polar degree
Proof that the latter remark computes the preimages under the gradient map is the same as for Theorem 5.40(4). The equivalence to the (usual) Newton number is implied by Lemma 5.33.
Figure 5: The polytope $P$ and the triangle $\Delta^{\circ}$ in the cone $C_{d}$
Recall Huh’s formula for the degree of the gradient map [Huh13, Corollary 4.6]. Consider a homogeneous degree $d$ polynomial $f\in\mathbb{C}[z_{0},\dots,z_{n}]$. Denote by $P\subset d\Delta^{\circ}$ the Newton polytope of $f$. Define
$m_{i}(P):=\operatorname{MV}_{n}(\underbrace{\Delta^{\circ},\ldots,\Delta^{\circ}}_{n-i},\underbrace{P,\ldots,P}_{i})\quad\text{for}\ i=0,\ldots,n.$
@@
([Huh13, Corollary 4.6]) Suppose that $f$ is nondegenerate with respect to its Newton polytope. Then the degree of the gradient map $\text{grad}(f):\mathbb{C}\mathbb{P}^{n}\to\mathbb{C}\mathbb{P}^{n}$ is the sum
@@
Denote by $E_{0}^{d},\dots,E_{n}^{d}$ the facets of $C_{d}$. Denote by $D_{i}$ the polytope $\operatorname{Conv}((\mathbf{P}\ast\mathbf{\Delta}^{\circ})\setminus E^{d}_{i})\subset C_{d}$. Then for a generic homogeneous degree $d$ polynomial with support $\mathbf{P}\subset\mathbb{Z}^{n+1}$, we have the following formula for the polar degree:
$\text{Pdeg}(f)=\operatorname{MV}(D_{0},\dots,D_{n})$
@@
The most interesting cases are the cases of the polar degree equals zero (vanishing hessian hypersurfaces) or one (homaloidal hypersurfaces corresponding to Cremona transformations); see e.g. [CRS08]. Note that due to the Proposition 5.51, the Propositions 5.6 and 5.7 provide criteria for these cases (in Newton non-degenerate case).
Question 5.53.
Can we classify polytopes $P\subset d\Delta^{\circ}$, so that $\nu_{C_{d}}(P\ast\Delta^{\circ})$ is equal to 0 or $1$?
For instance, we have a conjecture concerning Newton non-degenerate homaloidal hypersurfaces. The conjecture holds for the examples [Huh13, below Example 4.9] and we did not find any counterexamples.
@@
Consider the abstract lattice polyhedral poset $\mathcal{P}^{\bullet}$ defined as the agglutination of the polytope $P\ast\Delta^{\circ}$ and the face poset of the unit $(n+1)$-simplex $\Delta^{\bullet}$ along the face poset of the $n$-simplex $\Delta^{\circ}$. If we ignore the lattice structure of the polytopes in $\mathcal{P}^{\bullet}$, we see that it is isomorphic to a (non-lattice) subdivision of $\operatorname{Conv}(O_{d},\Delta^{\circ},P)$ into $P\ast\Delta^{\circ}$ and $\operatorname{Conv}(O_{d},\Delta^{\circ})$, where $O_{d}$ is the apex of the cone $C_{d}$. Note that $\operatorname{Conv}(O_{d},\Delta^{\circ},P)$ is a (non-lattice) convenient polytope in $C_{d}$. This induces strong formal subdivision with boundary $\sigma:(\mathcal{P}^{\bullet},\mathcal{P}^{\bullet}_{B})\to\mathcal{C}^{\pm}$, where $\mathcal{C}$ is the face poset of the cone $C_{d}$ (or, equivalently, the face poset of $\Delta^{\circ}$).
@@
$\text{Pdeg}(f)=\nu_{{C_{d}}}(P\ast\Delta^{\circ})=\frac{1}{2}\ell^{*}_{\mathcal{C}}(\mathcal{P}^{\bullet}\sharp_{\mathcal{P}^{\bullet}_{B}}\sharp\mathcal{P}^{\bullet};1)$
@@
The poset $\mathcal{C}$ is boolean, so the local $h_{\mathcal{C}}^{*}$-polynomial is the alternating sum of the corresponding $h_{\mathcal{C}}^{*}$-polynomials. The evaluation of the $h_{\mathcal{C}}^{*}$-polynomial of polytopes at $t=1$ is the lattice volume. The corresponding alternating sum of volumes for $\Delta^{\bullet}$ equals 0, and for $\Delta^{\circ}\ast P$ equals $2\nu_{{C_{d}}}(P\ast\Delta^{\circ})$. ∎
@@
$\text{Pdeg}(f)\leq\ell^{*}(P\ast\Delta^{\circ};1)$
6 $\mathbf{B_{k}}$-polytopes